OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-6, 1) m = 6

OpenStudy (anonymous):

maybe i should explain better so you can do these on your own (-6,1), this is a coordinate where when x=-6, y=1 m is your slope literally plug all of the values in, that is -6 for x, 1 for y and 6 for m y=mx+b 1=(6)(-6)+b b=1-(6)(-6) b=37 now use your slope and the value you find for b to create your equation y=6x+37